itzlanalol:

“Mask” Behind those wonderful eyes, The beautiful smirk, Ocean kindness, One amazing person, Has proven to Lana everything, He has showed her love, Kindness, Respect, And love, She loves him, Not because of the mask he wears, But because of who he is, He showed her only his ways, A true person would always show, Happiness, With his touch of love, Lana will never ever lose him, The way she did with her father, He didn’t show his mask like, Jack did for Lana, Respect, Love, And happiness, Is the only thing Lana wanted from Jack.

Swifty:

1 attachment
Swifty:

I like it

BriannaFloyd:

🔥🔥 looks nice

itzlanalol:

thank you

BriannaFloyd:

@itzlanalol wrote:
thank you
np

LavenderLuvBug:

This is beautifully written. I love the way you worded everything.

cutebunny123:

Reading this made my day! I'm looking forward to reading more of your stuff!

kai6665:

good job! 🔥🔥

AiannaFLOYD:

Dats amazing keep it up.

Sharkattack123:

amazing

itzlanalol:

THANK YOU

nunyaaaaaa:

its amazing i likee ittt :)

itzlanalol:

thanks

nunyaaaaaa:

no problem

itzlanalol:

good day

AiannaFLOYD:

No problem! Have a good day luv

velmalovesshaggy145:

yoz

velmalovesshaggy145:

jk whats up lan

alexyshull:

Wow that was amazing u did amazing keep it up!!!! :)

velmalovesshaggy145:

I love it

itzlanalol:

IT WOULD BE BUT THIS GIRL DOWN STOLE IT FROM ME

Sailor:

Nice job.

cynthiaclark:

its amazinggg

7469838583:

this was beautiful

pennalopee:

they are not even the same bc the parts cut off i didnt steal it from you True Love Behind those wonderful eyes, The beautiful smirk, His kindness, One amazing person, Has proven to her everything, He has showed not just her love but, Kindness, Respect, attention And love, She loves him, so much Not because of the things he does, But because of who he is, He showed her only his ways, Of a true person would always show, Happiness, With his touch of love,and caring.and never letting her down or doing anything to hurt her being there for her making her smile

itzlanalol:

@pennalopee wrote:
they are not even the same bc the parts cut off i didnt steal it from you True Love Behind those wonderful eyes, The beautiful smirk, His kindness, One amazing person, Has proven to her everything, He has showed not just her love but, Kindness, Respect, attention And love, She loves him, so much Not because of the things he does, But because of who he is, He showed her only his ways, Of a true person would always show, Happiness, With his touch of love,and caring.and never letting her down or doing anything to hurt her being there for her making her smile

pennalopee:

@itzlanalol wrote:
@pennalopee wrote:
they are not even the same bc the parts cut off i didnt steal it from you True Love Behind those wonderful eyes, The beautiful smirk, His kindness, One amazing person, Has proven to her everything, He has showed not just her love but, Kindness, Respect, attention And love, She loves him, so much Not because of the things he does, But because of who he is, He showed her only his ways, Of a true person would always show, Happiness, With his touch of love,and caring.and never letting her down or doing anything to hurt her being there for her making her smile
your saying i stole it from you

xxDeppressionxx:

$$\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}$$

pennalopee:

@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something

Scorpion13:

Cool

itzlanalol:

yee

itzlanalol:

@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK

pennalopee:

@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE

pennalopee:

@pennalopee wrote:
@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE
AND WHY I SAY SORRY WHEN I DIDNT STEAL YOU SHII

itzlanalol:

@pennalopee wrote:
@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE
omg they do you STEALER, you literally changed up the wording. IF YOU READ MY POEM YOU'D SEE IT YOURSELF LAIR.

pennalopee:

@itzlanalol wrote:
@pennalopee wrote:
@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE
omg they do you STEALER, you literally changed up the wording. IF YOU READ MY POEM YOU'D SEE IT YOURSELF LAIR.
BRUH IM NOT EVEN LYING BUT SAY WTV YOU WANT BC IK I DIDNT DO ANYTHING WRONG

itzlanalol:

@pennalopee wrote:
@itzlanalol wrote:
RIGHT LETS SEE WHAT THE MODS THINK THEN STEALER
@pennalopee wrote:
@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE
omg they do you STEALER, you literally changed up the wording. IF YOU READ MY POEM YOU'D SEE IT YOURSELF LAIR.
BRUH IM NOT EVEN LYING BUT SAY WTV YOU WANT BC IK I DIDNT DO ANYTHING WRONG

pennalopee:

@itzlanalol wrote:
@pennalopee wrote:
@itzlanalol wrote:
RIGHT LETS SEE WHAT THE MODS THINK THEN STEALER
@pennalopee wrote:
@itzlanalol wrote:
@pennalopee wrote:
@xxdeppressionxx wrote:
\[\displaystyle{\displaylines{\Large \color{Orange}{please~don’t~fight.}}}\]
we wont fight it was just two people misunderstanding something
right misunderstanding the fact you STOLE MY POEM AND WON'T SAY SORRY MHMMM OK
JUST LOOK AT THE DAMN THING YOUR BLIND THEY NOT REALLY THE SAME DUDE
omg they do you STEALER, you literally changed up the wording. IF YOU READ MY POEM YOU'D SEE IT YOURSELF LAIR.
BRUH IM NOT EVEN LYING BUT SAY WTV YOU WANT BC IK I DIDNT DO ANYTHING WRONG

XxBluexX:

amazing i love it <3

CrackedNuts:

I love this!

Olive2006:

I love this and it really makes people want to read it.

7469838583:

wow um this really changed alot lmao...

OLIVER69:

I love it. < 3

Flowerpower5290:

nice!!!!!!!!!